# Drops of volatile binary mixtures on brush-covered substrates

### v1

@@ Appendix A Free energy contributions
Here, we briefly discuss the individual contributions to the energy density, i.e., to the integrand of the functional (2) of the main text. First, we consider the brush energy per substrate area $f_{\mathrm{b}}$. As the brush consists of end-tethered polymer chains without cross-coupling, we use an amended Flory-Huggins free energy of mixing. In particular, as compared to a standard model of a ternary mixture the translational entropy of the polymers is replaced by an entropic spring term ultimately employing an Alexander-de Gennes approach [23, 24], as discussed in Ref. [19, 25]. Writing the expression in terms of $\zeta$’s we have
@@
where $\Lambda$ is the standard thermal de Broglie wavelength.
Considering a macroscopically thick flat layer of a pure liquid 1 or 2 in coexistence with saturated vapor, the influence of wetting and interface energies on the equilibrium state can be neglected. Then, the condition of equal chemical potentials reduces to
@@
brush state drop state
pure liquid 1 ($\phi_{\text{d1}}=1$)
pure liquid 2 ($\phi_{\text{d2}}=1$)
dry brush $(\phi_{\mathrm{bp}}=1)$
$\gamma_{\text{pl1}}$
$\gamma_{\text{pl2}}$
saturated with liquid 1 $(\phi_{\text{b1}}=1)$
0
$\gamma_{12}\ll\gamma_{\text{pl2}}$
saturated with liquid 2 $(\phi_{\text{b2}}=1)$
$\gamma_{12}\ll\gamma_{\text{pl1}}$
0
@@
$f_{\mathrm{wet}}(h_{\text{d}})=\frac{A}{2h_{\text{d}}^{2}}\,\left(\frac{2h_{\text{a}}^{3}}{5h_{\text{d}}^{3}}-1\right),$ (11)

### v2

@@ Appendix A Free energy contributions
Here, we briefly discuss the individual contributions to the energy density, i.e., to the integrand of the energy functional $\mathcal{F}$, i.e., Eq. (2) of the main text. First, we consider the brush energy per substrate area $f_{\mathrm{b}}$. As the brush consists of end-tethered polymer chains without cross-coupling, we use an amended Flory-Huggins free energy of mixing. In particular, as compared to a standard model of a ternary mixture the translational entropy of the polymers is replaced by an entropic spring term ultimately employing an Alexander-de Gennes approach [23, 24], as discussed in Ref. [19, 25]. Writing the expression in terms of $\zeta$’s we have
@@
where $\Lambda$ is the standard thermal de Broglie wavelength. Considering a macroscopically thick flat layer of a pure liquid 1 or 2 in coexistence with saturated vapor, the influence of wetting and interface energies on the equilibrium state can be neglected. Then, the condition of equal chemical potentials reduces to
@@
brush state drop state pure liquid 1 ($\phi_{\text{d1}}=1$) pure liquid 2 ($\phi_{\text{d2}}=1$)
dry brush $(\phi_{\mathrm{bp}}=1)$ $\gamma_{\text{pl1}}$ $\gamma_{\text{pl2}}$
saturated with liquid 1 $(\phi_{\text{b1}}=1)$ 0 $\gamma_{12}\ll\gamma_{\text{pl2}}$
saturated with liquid 2 $(\phi_{\text{b2}}=1)$ $\gamma_{12}\ll\gamma_{\text{pl1}}$ 0
@@
$f_{\mathrm{wet}}(h_{\text{d}})=\frac{A}{2h_{\text{d}}^{2}}\,\left(\frac{2h_{\text{a}}^{3}}{5h_{\text{d}}^{3}}-1\right)$ (11)